OpenStudy (anonymous):

Solve: -1/2 x + 1 = −x + 8

OpenStudy (anonymous):

minus 1 add x you have your answer

OpenStudy (anonymous):

so?

OpenStudy (anonymous):

x=14

OpenStudy (anonymous):

1x/2+x=8-1

OpenStudy (anonymous):

-1x/2+2x/2=7

OpenStudy (anonymous):

x/2=7

OpenStudy (anonymous):

x=14